OpenStudy (anonymous):

definite integral of x+3 lower limit -1 upper limit 3. Step by step please

OpenStudy (unklerhaukus):

hi \[\int\limits_{-1}^3(x+3)\;\mathrm dx =\frac{x^{1+1}}{1+1}+3\frac{x^{0+1}}{0+1}\Big|_{-1}^3\]

OpenStudy (unklerhaukus):

does this step make sense,? each term in the integrand has had its index of x raised by one, and then that term is divided by the new index

OpenStudy (anonymous):

Hello, yes so far it makes sense.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

so now simplify those denominators and indexes , what do you get ?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(x^2+3)\]

OpenStudy (anonymous):

opps 3x

OpenStudy (unklerhaukus):

check the denominator of that second term

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(x^2+3x)\]

OpenStudy (unklerhaukus):

almost, but not quite

OpenStudy (unklerhaukus):

1+0=1≠2

OpenStudy (anonymous):

\[\frac{ x^2 }{ 2 }+3x\]

OpenStudy (unklerhaukus):

that is better! so you have \[\int\limits_{-1}^3(x+3)\;\mathrm dx =\frac12x^2+3x\Big|_{-1}^3\] now to evaluate the limits \[=\Big(\tfrac12(3)^2+3(3)\Big)-\Big(\tfrac12(-1)^2+3(-1)\Big)\]

OpenStudy (anonymous):

So the final answer will be 16?

OpenStudy (unklerhaukus):

\[\large\color{red}\checkmark\]

OpenStudy (anonymous):

Thanks!(: